# DeepCompile: A Compiler-Driven Approach to Optimizing Distributed Deep Learning Training

### v1

@@ 3. System Design
Figure 2 shows the overall workflow. Starting from a user-defined training script written in a framework such as PyTorch, a base compiler (e.g., the PyTorch compiler) lowers the model into an intermediate representation (IR) as a computation graph. DeepCompile takes this graph as input, transforms it by adding communication operations for distributed training, and applies optimizations. The resulting graph is then deployed to a runtime engine running on GPU servers.
Figure 2. Workflow of compilation and optimization with DeepCompile
@@
To support dynamic and memory-aware optimizations, DeepCompile organizes its graph transformations as a sequence of optimization passes. Each pass rewrites the computation graph based on a specific strategy, such as inserting communication operators or reordering memory-intensive computations. After applying each pass, DeepCompile executes the modified graph to collect runtime profiling data, including operator execution times, communication overhead, and memory usage trends. This information is then used to guide the next optimization pass. By repeating this process, DeepCompile incrementally refines the graph with awareness of runtime behavior.
@@
Figure 3 illustrates this two-level loop: an inner loop of optimization and profiling, and an outer loop that periodically runs training to reflect changes in the runtime environment. By alternating between these loops, DeepCompile adapts its transformations to the evolving memory and execution characteristics of the model. This enables coordinated application of optimizations such as prefetching, unsharding, and offloading in a unified and data-driven manner.
Figure 3. Optimization and profiling loop in DeepCompile
@@ 4. Optimizations
DeepCompile enables a broad range of optimizations by leveraging a computational graph extracted by the base compiler. In this paper, we first demonstrate how to implement the fully-sharded approach. We then describe additional optimizations to reduce communication overhead and reduce memory requirements.
@@ 4.1. Fully-sharded approach
Existing systems such as ZeRO-3 and FSDP insert all-gather and memory release operations around each layer that uses a given parameter. When a model has many layers with small parameter tensors, this results in frequent all-gather operations on small fragments. Since all-gather communication is inefficient for small data sizes (Fig. 5(c)), this leads to significant communication overhead. To mitigate this, both ZeRO-3 and FSDP allow users to manually group multiple layers into a larger block, reducing the number of all-gather operations. However, this increases the lifetime of the buffers holding gathered parameters. Furthermore, none of the layers in the block can begin computation until all of their associated all-gather operations have completed, often resulting in significant idle time. As a result, manually tuning the block size for optimal performance is difficult.
DeepCompile operates by modifying the computation graph to insert all-gather and release operations in a similar manner to existing frameworks. However, unlike systems that rely on layer boundaries or manually defined blocks, DeepCompile uses operator-level dependency analysis to determine the optimal placement of these operations.
@@
Figure 4. Graph modification
Although this scheduling improves memory efficiency, it does not necessarily improve execution speed. DeepCompile addresses this by applying a series of optimization passes. The initial pass focuses on ensuring the model fits within available memory, and subsequent passes improve training efficiency by adjusting the timing of communication, avoiding unnecessary sharding, and overlapping data transfers with computation.
@@ 4.2. Proactive prefetching
The initial graph transformation for the fully sharded approach in DeepCompile produces behavior similar to that of ZeRO-3 or FSDP operating in their most fine-grained configuration, where each layer triggers an all-gather before its computation and a release immediately after. However, because DeepCompile expresses this logic at the level of the computation graph, it enables detailed analysis of operator dependencies and memory usage patterns. This graph-based representation provides a foundation for more precise and globally coordinated optimizations, such as prefetching, which go beyond what is possible with heuristic or manually configured approaches.
With the fully sharded approach implemented in the computation graph, DeepCompile can build on top of it by introducing prefetching. Prefetching is a commonly used technique in fully sharded training systems, and both ZeRO-3 and FSDP include built-in support for it. The goal is to improve training efficiency by overlapping communication with computation, reducing the impact of all-gather latency. Figure 5(a) illustrates the timeline of a naive behavior of fully sharded approach, where green boxes represent computation and blue boxes indicate all-gather communication. The number in each blue box corresponds to the respective layer. Since all-gather communication is the primary source of overhead, both ZeRO-3 and FSDP employ prefetching to mitigate its impact. By launching all-gather earlier than the parameters are used, these methods increase opportunities to overlap communication and computation (Fig. 5(b)).
Figure 5. Parameter sharding and prefetching
In existing frameworks, prefetching is implemented in a static and coarse-grained manner. ZeRO-3 issues prefetches based on a predefined buffer size, launching all-gather operations for as many upcoming parameters as can fit in that buffer. Additionally, by grouping multiple layers into a larger block, ZeRO-3 and FSDP can fuse several all-gather operations. This improves communication efficiency, particularly for small data sizes. However, as discussed earlier, it can also delay the computation of earlier layers because the system must wait until all parameters involved in the fused communication are gathered.
Furthermore, ZeRO-3 and FSDP are not designed to account for dynamic changes in memory availability. As shown in Figure 1, memory usage fluctuates throughout the forward and backward passes. More memory tends to be available in the early stages of the forward pass and the later stages of the backward pass. However, existing methods do not effectively take advantage of these patterns. As a result, prefetching opportunities are often underutilized, leading to suboptimal efficiency.
@@
To address this limitation, we propose an optimization pass for proactive prefetching. Figure 6 illustrates the memory usage trend and scheduling behavior with and without proactive prefetching. The goal of this optimization pass is to move all-gather operations as early as possible while ensuring that memory constraints are not violated. Since the memory usage of each operator and the size of the all-gather buffers are known from profiling, it is possible to estimate the total memory footprint if an all-gather is scheduled earlier than usual.
@@
In Fig. 6, the dotted line shows the effect of proactive prefetching. In this example, there are three all-gather operations. Based on profiling results, the system determines that prefetching the first two can be safely done at the beginning without exceeding the memory limit. The third all-gather, however, would exceed the memory budget if issued immediately. Therefore, it is scheduled only after the buffer from the first parameter has been released. This scheduling strategy ensures efficient memory usage and allows prefetching to be applied as aggressively as possible within the memory budget.
Algorithm 1 formalizes this procedure. It examines each all-gather operation and determines whether it can be moved earlier in the schedule without exceeding the memory limit, based on the profiled memory usage and buffer size. The algorithm iterates over the operators in the initial schedule $S_{0}$ in reverse order. If the operator $o_{i}$ is an all-gather, it checks whether the operator can be moved earlier without violating the memory constraint.
@@
$N$ Number of operators
$S_{0}$ Initial schedule of operators $[o_{1},o_{2},\ldots,o_{N}]$
$OS$ Fragments of optimizer states $[os_{1},os_{2},\ldots]$
$\text{P}_{\text{mem}}(o)$ Profiled memory usage just before operator $o$
@@
1:  $S\leftarrow[]$ // Operators in new schedule
2:  $U\leftarrow[]$ // Unscheduled all-gathers
3:  for $i=N$ to $2$ do
@@
If the estimated memory usage remains below the limit, the all-gather operator $o_{i}$ is added to a group $U$ of unscheduled all-gather operators that can be moved earlier. If the memory usage exceeds the limit, the algorithm schedules the all-gather operators in $U$ at the current position, fusing them where appropriate using the Fuse function. This strategy ensures that memory usage stays within the limit $M$.
We also set the limit $M_{\text{prefetch}}$ on the prefetch group, as moving too many allgather operations earlier does not provide any additional benefit. We can set this limit to a significantly larger value than the predefined prefetch buffer size used in existing implementations such as ZeRO-3 because we also ensure the total memory usage is less than the limit $M$. By imposing this limit, we ensure that some memory remains available for selective unsharding, which is discussed in the next section.
When the data size is small, the communication time remains nearly constant, even as the data size increases. Therefore, fusing all-gather operators reduces communication overhead (Fig. 5(c)(d)), especially for small data sizes. The criterion for fusing two all-gather operations is as follows: if $T_{c}(V_{1})+T_{c}(V_{2})>\alpha\cdot T_{c}(V_{1}+V_{2})$, where $V_{1}$ and $V_{2}$ are the data sizes of two operations, fusion is beneficial and the operations are merged. Otherwise, they are left separate. Here, $T_{c}(V)$ is profiled at runtime, and $\alpha$ is a tunable parameter, set to 1.5 in our experiments.
SimpleFSDP (Zhang et al., 2024) takes a similar approach to ours by optimizing prefetching for the fully sharded approach using a compiled computation graph. It performs memory-aware fusion of all-gather operations to reduce communication overhead. However, SimpleFSDP does not consider how memory usage dynamically changes during the forward and backward passes to determine how early each all-gather can be issued. In contrast, our proactive prefetching algorithm uses profiling data to make this decision, allowing communication to be overlapped with computation while staying within memory constraints.
@@ 4.3. Selective unsharding
After applying the optimization pass for proactive prefetching, some free memory may remain, since the prefetching size is deliberately limited. Selective unsharding leverages this remaining memory by keeping some parameters unsharded after they are gathered. This approach can significantly reduce communication overhead, although it requires additional memory.
In the fully sharded approach, parameters are distributed across all GPUs, and all-gather operations are required after each parameter update. Selective unsharding, however, allows gathered parameters to remain unsharded after the forward pass, eliminating the need for all-gather in the backward pass.
This technique is especially beneficial when used with gradient accumulation, a common method to increase the effective batch size without increasing per-GPU memory usage. Given a gradient accumulation step of $N$, the model performs $N$ forward and backward passes while accumulating gradients before applying a single parameter update. Since parameters are not updated during this period, they can remain unsharded across multiple forward and backward passes.
We run profiling after applying proactive prefetching to measure the peak memory usage. Then, we select as many parameters to keep unsharded as possible, ensuring that the total buffer size does not exceed the memory limit. The selection considers both the buffer size of each all-gather operator $o$, denoted as $\text{B}(o)$, and its communication time $T_{c}(\text{B}(o))$. Parameters are prioritized based on the ratio $\frac{T_{c}(\text{B}(o))}{\text{B}(o)}$, where a higher value indicates a greater reduction in communication time relative to memory cost. Following the trend shown in Fig. 5(d), smaller parameters are generally selected first.
@@ 4.4. Adaptive offloading
Although optimizer states such as momentum and variance used in Adam optimizer (Kingma and Ba, 2014) are only needed for parameter updates and are not required during the forward and backward passes, they occupy a significant amount of memory throughout training. Existing frameworks, such as DeepSpeed, provide functionality to offload these states to CPU memory. In fully sharded approaches like ZeRO-3, both model parameters and optimizer states are partitioned across GPUs. This design reduces the memory footprint on each device and enables parallel data transfers between GPUs and CPU memory. As a result, optimizer state offloading is often used together with fully sharded training.
@@
0:  Initial schedule $S_{0}$, Memory limit $M$, optimizer state fragments $OS$
@@
However, offloading and reloading optimizer states can typically significant overhead. To address this challenge, we propose an approach called adaptive offloading, which improves upon existing offloading mechanisms in two ways. First, by monitoring actual memory usage during training, it minimizes the amount of data that needs to be offloaded. Only the portion of optimizer states that would exceed the memory limit is transferred to the host. Second, it takes advantage of the characteristic memory usage patterns in training. Since memory usage tends to increase during the forward pass and decrease during the backward pass, adaptive offloading schedules data transfers to overlap with computation. This overlap reduces the performance impact of moving data between GPU and CPU memory.
Algorithm 2 presents the pseudo-code for the forward pass with adaptive offloading of optimizer states. The optimizer states are divided into a large enough number of fragments. At the start of the forward pass, the algorithm schedules operators to initiate asynchronous offloading of fragments that exceed the memory limit. Since the copy operation is asynchronous, computation can run in parallel. The memory must only be freed after the copy operation completes. Therefore, the peak memory usage before each operator is checked based on profiling when scheduling. If the memory usage exceeds the limit, the algorithm adds operators to synchronize the copy and free memory for the fragments.
We also iterate over the operators in the backward pass schedule, checking the available memory. If sufficient memory is available from that operator to the end, we schedule a reload operation to transfer the optimizer states back to the GPU. This approach enables overlapping the backward computation with the data transfer from host memory to GPU memory while adhering to memory constraints. Since memory usage often increases during forward passes and decreases during backward passes, this strategy allows for efficient overlap of computation and offloading/reloading operations.
@@ 5.1. Experimental settings
The experiments were performed on two or four servers, each equipped with eight NVIDIA H100 GPUs (80GB memory per GPU) and 1TB of system memory. The GPUs within each node were interconnected via NVLink, and each server included eight network interfaces (HCAs) connected via InfiniBand. The nodes were interconnected using InfiniBand with an observed bandwidth of 340GB/s for both allgather and reduce-scatter operations. Each server was equipped with two Intel(R) Xeon(R) Platinum 8462Y+ processors.
In our experiments, we used the Llama-3 70B and Mixtral 8x7B models (47 billion parameters)(Jiang et al., 2024) as our target workloads. These models are representative examples of dense and Mixture-of-Experts (MoE) architectures(Shazeer et al., 2017; Lepikhin et al., 2021; Fedus et al., 2022), respectively. Both were run using bfloat16 with mixed precision (Micikevicius et al., 2017), while maintaining FP32 copies of parameters, gradients, and Adam optimizer states. We also employed activation checkpointing, integrated into the model implementation at the Transformer layer level, where each layer is recomputed as a block.
We used PyTorch v2.6.0 and models from the HuggingFace Model Hub without any built-in parallelization support. The experiments were conducted with Python 3.10 and CUDA 12.4. DeepSpeed v1.6.4 and PyTorch FSDP were used as the baseline frameworks, and DeepCompile was built on top of DeepSpeed. The optimization passes and profilers in DeepCompile were implemented in Python (approximately 2.6K lines), while the custom operators injected into the computational graph were implemented in C++ (approximately 900 lines).
While some distributed training strategies support partial sharding—for example, sharding only optimizer states, or both gradients and optimizer states while keeping parameters unsharded—these configurations are not sufficient for our target models. The total parameter sizes of Llama-3 70B and Mixtral 8x7B exceed the memory capacity of a single GPU (140 GB and 94 GB, respectively), requiring all parameters, gradients, and optimizer states to be fully sharded across devices. For this reason, we use the fully sharded configurations of both DeepSpeed ZeRO-3 and PyTorch FSDP in our evaluation.
Other large-scale training frameworks, such as Megatron-LM, are not included in our comparison. Although Megatron-LM has been widely adopted for large model training, it requires a specialized implementation where parallelism strategies are tightly coupled with model code. In contrast, we focus on model-agnostic approaches that can be applied to standard implementations. Similarly, frameworks that automate parallelization, such as Alpa (Mo et al., 2022), FlexFlow (Jia et al., 2018), and Unity (Unger et al., 2022), are not included, as they remain experimental and do not currently support the models used in this evaluation.

### v2

@@ 3. System Design
Figure 2 shows the overall workflow. Starting from a user-defined training script and a model implementation written in a framework such as PyTorch, a base compiler (e.g., the PyTorch compiler) lowers the model into an intermediate representation (IR) as a computation graph. DeepCompile takes this graph as input, transforms it by adding communication operations for distributed training, and applies optimizations. The resulting graph is then deployed to a runtime engine running on GPU servers.
Figure 2. Workflow of DeepCompile. DeepCompile transforms the IR generated by a base compiler, injecting distributed training optimizations before execution on GPUs.
@@
To support dynamic and memory-aware optimizations, DeepCompile organizes its graph transformations as a sequence of optimization passes, as is common in existing deep learning compilers. Each pass rewrites the computation graph based on a specific strategy, such as inserting communication operators or reordering memory-intensive computations. After applying each pass, DeepCompile executes the modified graph to collect runtime profiling data, including operator execution times, communication overhead, and memory usage trends. This information is then used to guide the next optimization pass. By repeating this process, DeepCompile incrementally refines the graph with awareness of runtime behavior.
@@
Figure 3 illustrates this two-level loop: an inner loop of optimization and profiling for each optimization pass, and an outer loop that periodically runs training to reflect changes in the runtime environment. By alternating between these loops, DeepCompile adapts its transformations to the evolving memory and execution characteristics of the model. This enables the coordination of multiple optimizations such as prefetching, unsharding, and offloading in a unified and profile-guided manner.
Figure 3. Profiling-guided optimization loop in DeepCompile. DeepCompile applies optimizations in iterative phases, with periodic training to reflect memory dynamics and coordinate passes such as prefetching and offloading.
@@ 4. Optimizations
DeepCompile enables a broad range of optimizations by leveraging a computational graph extracted by the base compiler. In this paper, we begin by describing how the fully sharded approach is expressed within our framework, and then introduce additional optimizations that reduce communication overhead and memory requirements.
@@ 4.1. Fully-sharded approach
Figure 4. Inserting all-gather and release operations for the fully sharded approach. This minimizes buffer lifetime by aligning communication and memory operations with actual usage.
In existing implementations of the fully sharded approach, such as ZeRO-3 and FSDP, communication and memory operations are typically triggered by hooks inserted at layer boundaries in the Python module structure. These hooks invoke all-gather operations and manage memory allocation and release in response to the forward and backward pass.
In contrast, DeepCompile operates by modifying the computation graph to insert all-gather and release operations, similar to existing frameworks. However, unlike systems that rely on layer boundaries or manually defined blocks, DeepCompile can flexibly inject or reorder such communication and memory operations based on the dependencies among operators.
@@
Although this scheduling improves memory efficiency, it does not necessarily improve execution speed. To address this, DeepCompile applies a series of optimization passes. The initial pass ensures that the model fits within available memory, while subsequent passes improve training efficiency by adjusting the timing of communication, avoiding unnecessary sharding, and overlapping data transfers with computation.
@@ 4.2. Proactive prefetching
The initial graph transformation for the fully sharded approach in DeepCompile produces behavior similar to that of ZeRO-3 and FSDP, where each layer triggers an all-gather before its computation and releases the gathered parameters immediately afterward. Figure 5(a) illustrates this baseline behavior during a forward pass. Computation (green boxes) and all-gather communication (blue boxes) are executed sequentially per layer, resulting in no overlap.
Figure 5. Impact of prefetching on memory usage and communication–computation overlap. (a) Without prefetching, all-gather operations are issued just before computation, resulting in no overlap. (b) Prefetching enables overlap, but fixed buffer sizes may cause out-of-memory errors. A dynamic strategy adapts to memory availability and avoids such failures.
The bottom track in each timeline shows memory usage over time. Activation memory (green), used for intermediate computation, accumulates across layers and is retained until the backward pass. In contrast, memory used for all-gather communication (blue) increases when the operation begins and is released once the corresponding computation completes. As a result, more memory is typically available earlier in the forward pass and later in the backward pass.
To reduce communication overhead, systems such as ZeRO-3 and FSDP support prefetching, which launches all-gather operations earlier than strictly required to enable communication–computation overlap. Figure 5(b) demonstrates how different prefetching strategies impact memory usage and execution behavior. With a fixed prefetching buffer size of 2, safe overlap is achieved without exceeding memory limits. However, a larger buffer size of 4 causes an out-of-memory error due to increased memory pressure from overlapping communication and accumulating activation memory.
To prevent such failures, ZeRO-3 and FSDP employ a static buffer size and limits prefetching to the parameters that fit within that bound. While safe, this conservative strategy fails to adapt to dynamic changes in memory availability during training. As illustrated in Figure 1, memory usage fluctuates significantly throughout execution. Fine-grained control is required to determine exactly when to initiate each all-gather operation. The bottom timeline in Figure 5(b) exemplifies such dynamic scheduling, which is not feasible in existing systems like ZeRO-3 or FSDP.
@@
To address this limitation, we introduce an optimization pass for proactive prefetching, which schedules all-gather operations as early as possible while ensuring that memory constraints are not violated. By profiling memory usage and all-gather buffer sizes after applying the optimization pass for the fully sharded approach, DeepCompile can estimate the total memory footprint of earlier communication.
@@
Figure 6 illustrates how this scheduling works. Solid lines indicate memory usage without prefetching, while dotted lines show usage with proactive prefetching. In this example, three all-gather operations are scheduled. Based on the memory budget, the first two can be safely prefetched at the beginning. The third, however, would exceed the memory limit if issued immediately, so it is delayed until after the buffer for the first parameter has been released. This approach maximizes prefetching opportunities while maintaining safe memory usage throughout execution.
@@
NN Number of operators
OSOS Fragments of optimizer states $os_{1},os_{2},\ldots]$
$\text{P}_{\text{mem}}(o)$ Profiled memory usage before operator $o$
@@
[…]
2:U[$\leftarrow[] // U$nscheduled all-gathers
3:for $i=N$ to 22 do
@@
Algorithm 1 formalizes this procedure. It iterates over the operators in the initial schedule $S_{0}$ in reverse order. The schedule $S_{0}$ is produced by the preceding optimization pass that enables the fully sharded approach. For each operator $o_{i}$, if it is an all-gather operation, the algorithm checks whether it can be moved earlier in the schedule without violating the memory constraint. This decision is based on the profiled memory usage and the required buffer size.
If the estimated memory usage remains below the limit, the all-gather operator $o_{i}$ is added to a group $U$ of unscheduled all-gather operators that can be moved earlier. If the usage exceeds the limit, the algorithm schedules the operators in $U$ at the current position, applying fusion based on the Fuse function. This strategy maintains memory usage within the limit $M$.
We also set the limit $M_{\text{prefetch}}$ on the prefetch group, as moving too many all-gather operations earlier does not provide any additional benefit. By imposing this limit, we ensure that some memory remains available for selective unsharding, which is discussed in the next section.
The Fuse function decides whether to fuse a set of all-gather operators and returns a sequence of fused or unfused operations. For small message sizes, communication time increases slowly and remains nearly flat. As a result, combining multiple small all-gather operations into a single larger one can significantly reduce communication overhead. The $\text{Fuse}(U)$ operation fuses two all-gather calls if the following condition holds: if $T_{c}(V_{1})+T_{c}(V_{2})>\alpha\cdot T_{c}(V_{1}+V_{2})$, where $V_{1}$ and $V_{2}$ are the data sizes of two operations. Here, $T_{c}(V)$ is profiled communication time for size $V$, and $\alpha$ is a tunable parameter.
SimpleFSDP (Zhang et al., 2024) takes a similar approach to ours to optimize the fully sharded approach using a compiled computation graph. It performs memory-aware fusion of all-gather operations to reduce communication overhead. However, SimpleFSDP does not consider how memory usage dynamically changes during the forward and backward passes to determine how early each all-gather can be issued. In contrast, our proactive prefetching algorithm uses profiling data to make this decision, allowing communication to be overlapped with computation while staying within memory constraints.
@@ 4.3. Selective unsharding
After applying the optimization pass for proactive prefetching, some free memory may remain. Selective unsharding leverages this remaining memory by keeping certain parameters in their unsharded form after they are gathered and delaying their release until the parameter update step, which occurs after the backward pass. By avoiding repeated all-gather operations for these parameters, this approach reduces communication overhead.
This technique is especially beneficial when used with gradient accumulation, a common method to increase the effective batch size without increasing GPU memory usage. Given a gradient accumulation step of $n$, the model performs $n$ forward and backward passes while accumulating gradients before updating parameters. Since no parameter updates occur during this period, the gathered parameters can remain unsharded across multiple forward and backward passes. After a backward pass, the optimizer states are reloaded to GPU memory to update parameters.
We run profiling after applying proactive prefetching to measure the peak memory usage. Then, we select as many parameters to keep unsharded as possible, ensuring that the total buffer size does not exceed the memory limit. The selection considers both the buffer size of each all-gather operator $o$, denoted as $\text{B}_{\text{ag}}(o)$, and its communication time $T_{c}(\text{B}_{\text{ag}}(o))$. Parameters are prioritized based on the ratio $\frac{T_{c}(\text{B}_{\text{ag}}(o))}{\text{B}_{\text{ag}}(o)}$, where a higher value indicates a greater reduction in communication time relative to memory cost. Since communication is less efficient for smaller messages, parameters with smaller buffer sizes are generally selected first.
@@ 4.4. Adaptive offloading
Optimizer states such as momentum and variance used in the Adam optimizer (Kingma and Ba, 2014) require a lot of GPU memory, but they are only needed for parameter updates and not during the forward or backward pass. Therefore, offloading them to CPU memory can significantly reduce peak memory usage. At the end of each backward pass, the optimizer states must be loaded back into GPU memory for parameter updates. As implemented in ZeRO-3, the fully sharded approach is often combined with offloading, as both techniques help reduce memory requirements when training large-scale models.
@@
0: Initial schedule $S_{0}$, Memory limit $M$, optimizer state fragments $OS=[os_{1},os_{2},\ldots]$
@@
However, data transfers for offloading and reloading optimizer states can incur significant overhead. To address this challenge, we propose an approach called adaptive offloading, which improves upon existing offloading mechanisms in two ways. First, by monitoring actual memory usage during training, it minimizes the amount of data that needs to be offloaded. Only the portion of optimizer states that would exceed the memory limit is transferred to the host. Second, it takes advantage of the characteristic memory usage patterns in training. Since memory usage tends to increase during the forward pass and decrease during the backward pass, adaptive offloading schedules data transfers to overlap with computation. This overlap reduces the performance impact of moving data between GPU and CPU memory.
Algorithm 2 presents the pseudo-code for the forward pass with adaptive offloading of optimizer states. The optimizer states are divided into a large enough number of fragments. At the start of the forward pass, the algorithm schedules operators to initiate asynchronous offloading of fragments that exceed the memory limit. Since the copy operation is asynchronous, computation can run in parallel. The memory must only be freed after the copy operation completes. Therefore, the peak memory usage before each operator is checked based on profiling when scheduling. If the memory usage exceeds the limit, the algorithm adds operators to synchronize the copy and free memory for the fragments. As memory usage typically increases during the forward pass, this strategy enables efficient overlap of computation and offloading.
We also iterate over the operators in the backward pass schedule, checking the available memory. If sufficient memory is available from that operator to the end, we schedule an operator to initiate asynchronous transfer the optimizer states back to the GPU. This approach enables overlapping the backward computation with the data transfer from host memory to GPU memory while adhering to memory constraints. In contrast to the forward pass, memory usage decreases during the backward pass, making it easier to hide reloading overhead.
4.5. Composability
DeepCompile organizes optimizations as modular compiler passes, enabling users to compose them flexibly based on their training goals and constraints. However, the effectiveness of these passes depends not only on their individual design but also on their application order and compatibility.
The optimization passes for proactive prefetching, selective unsharding, and adaptive offloading are designed to be applied after the pass that enables the fully sharded approach, which rewrites the computation graph to insert all-gather and release operations for sharded parameters. However, the order and combination of these passes can significantly affect overall performance.
A beneficial application order is to first apply proactive prefetching, followed by selective unsharding. This ordering ensures that selective unsharding can make use of the remaining memory after prefetching has opportunistically advanced communication while staying within memory constraints. In contrast, applying selective unsharding before proactive prefetching can reduce the effectiveness of prefetching. Since unsharding attempts to utilize as much available memory as possible to retain parameters in memory, little capacity remains for prefetch buffers. As a result, the benefits of prefetching are largely lost.
Adaptive offloading is designed for scenarios where the model does not fit in GPU memory even after sharding. In practice, when adaptive offloading is applied, the memory available for proactive prefetching and selective unsharding is already limited, and the additional benefits from these passes tend to diminish.
@@ 5.1. Experimental settings
The experiments were performed on two or four servers, each with eight NVIDIA H100 GPUs (80GB per GPU), 1TB of system memory, and two Intel(R) Xeon(R) Platinum 8462Y+ CPUs. GPUs within each server were connected via NVLink, and servers were interconnected using InfiniBand (340GB/s observed bandwidth for allgather and reduce-scatter). Each server also included eight InfiniBand-connected HCAs.
We used Llama-3 70B and Mixtral 8x7B (47B) as representative dense and Mixture-of-Experts (MoE) models (Jiang et al., 2024; Shazeer et al., 2017; Lepikhin et al., 2021; Fedus et al., 2022). Both ran in bfloat16 with mixed precision (Micikevicius et al., 2017), keeping FP32 copies of parameters, gradients, and Adam states. We applied activation checkpointing (Chen et al., 2016) at the Transformer layer level, recomputing each layer as a block.
We used PyTorch v2.6.0 and models from the HuggingFace Model Hub without any built-in parallelization support. The experiments were conducted with Python 3.10 and CUDA 12.4. ZeRO-3 (DeepSpeed v1.6.4) and PyTorch FSDP were used as the baseline frameworks, and DeepCompile was built on top of DeepSpeed. The optimization passes and profilers in DeepCompile were implemented in Python (approximately 2.6K lines), while the custom operators injected into the computational graph were implemented in C++ (approximately 900 lines).
Partial sharding strategies (e.g., sharding only optimizer states) are insufficient for our target models. Llama-3 70B and Mixtral 8x7B exceed single-GPU memory (140 GB and 94 GB), requiring all parameters, gradients, and optimizer states to be fully sharded. Thus, we use the fully sharded configurations of ZeRO-3 and FSDP in our evaluation.
Other large-scale training frameworks, such as Megatron-LM, are not included in our comparison. Although Megatron-LM has been widely adopted for large model training, it requires a specialized implementation where parallelism strategies are tightly coupled with model code. In contrast, we focus on model-agnostic approaches that can be applied to standard implementations. Similarly, frameworks that automate parallelization, such as Alpa (Mo et al., 2022), FlexFlow (Jia et al., 2019), nnScalar (Lin et al., 2024), and Unity (Unger et al., 2022), are not included, as they do not currently support the models used in this evaluation.